# AlphaNet: Scaling Up Local Frame-based Neural Network Interatomic Potentials

### v3

@@ 2.3 AlphaNet is efficient—achieves the best speed-accuracy trade-off
The last efficiency evaluation is on the OC20 dataset, as shown in Table 5. Notably, EScAIP demonstrates remarkable efficiency in both speed and memory usage despite its large size, while AlphaNet shows a slight advantage in inference speed but consumes more memory than ut. It is important to note that models such as AlphaNet, SchNet, and DimeNet++ are conservative, meaning their force calculations inherently require additional computational resources for PyTorch autograd, resulting in higher inference time and memory costs.
@@ 3 Discussion
AlphaNet presents a significant advancement as a scalable neural network interatomic potential, offering both high precision and fast computational speed. This success can be attributed to its local-frame-based architecture, which scalarizes geometric quantities over local frames and effectively aggregate local frames to construct local- and global-aware features. By utilizing this architecture, our model excels at capturing complex interactions across a broad spectrum of systems, including metallic, covalent, and ionic bonding, as well as long-range interactions.
One of the key advantages of AlphaNet lies in its scalability. Specifically, our model is capable of maintaining high performance as both the size of the dataset and size of the atomic systems increase. Despite the growing complexity introduced by larger datasets, AlphaNet operates with fewer parameters, making it more efficient while still delivering robust results. This ability to handle large datasets without sacrificing accuracy positions AlphaNet as a promising solution for extensive simulations and real-world applications.
For future work, we aim to extend the model’s capabilities to explore additional types of interactions, such as hydrogen bonding, which are critical in a wide range of chemical and biological systems. Furthermore, the interpretability of the proposed model is worth exploring. While our current implementation achieves impressive results, gaining a deeper understanding of how the model interprets and processes interactions at the atomic level will enhance its applicability and trustworthiness in practical scenarios. There is another long-term challenge: the efficient use of GPU memory during training, particularly when dealing with large systems consisting of over 1,000 atoms (though we can parallelize it across multiple GPUs). This is a common limitation shared by most message-passing neural networks, and addressing it will be crucial for enabling the model to scale further in both size and complexity. Future work will focus on optimizing memory usage to overcome this bottleneck and ensure the model’s applicability to even larger systems.
@@ 4 Methodology
Equivariant message-passing neural networks. Equivariance is another crucial property to building MLFFs as molecular systems are invariant to the Euclidean group. A function $f$ is said to be equivariant with respect to certain group $G$ if $f\circ g(x)=g\circ f(x)$, $\forall x\in X,\,g\in G$. One natural property of message-passing neural networks is the permutation equivariance such that changing the order of the input set of nodes do not result in difference output for the same node as it only depends on the neighbors which are unchanged under permutation. In addition to permutation, we need to consider the special Euclidean group in 3D, SE(3), which includes rotation and translation operations. We do so by leveraging local frames (Du et al., 2022). Specifically, we build a set of equivariant and complete frames and scalarize the geometric quantity which does not break SE(3)-equivariance under nonlinear neural network encodings.
To build equivariant and complete frames from atomic positions, we consider the following edge-based frames following (Du et al., 2022):
@@
Efficient and expressive equivariant message-passing neural networks. Following (Du et al., 2024), we leverage an efficient and expressive extension of the above equivariant message-passing neural networks based on local frames. We consider two modules to improve the expressiveness while maintaining the efficiency of the network. Specifically, we leverage a local structure encoding module which scalarizes not only features from local neighbors but also from overlapping neighbors along each edge, denoted as $A_{ij}:=f_{l}(\text{Scalarize}(S_{i-j},\mathcal{F}_{ij}))$, where $f_{l}$ denotes a neural network and $S_{i-j}$ represents overlapping neighbors of node $i$ and $j$. In addition, we use an additional frame transition module which encodes the alignment between each neighboring frame to improve the expressiveness. We realize this by incorporating vector-based message passing to amortize the cost where $\mathbb{m}_{ij}$ denotes equivariant edge features. This framework is highly efficient as it avoids expensive higher-body message-passing neural networks and higher-order tensor updates.
@@
Rotary positional embedding as invariant frame transition. Positional embedding are known to improve and stabilize the training of transformers (Shaw et al., 2018). We introduce the elegant relationship between the commonly used Rotary Position Embedding (RoPE) and frame transition we use in our architecture (Su et al., 2024). In fact the frame transition can be considered as a generalized rotary positional embedding for 3D equivariant frames. Note that the spirit of rotary position embedding is to design an embedding scheme for relative position in sequential data. To build the rotary positional embedding,  (Su et al., 2024) utilizes the relative index between nodes $x_{i}$ and $x_{j}$ of a one-dimensional sequence to find a function $g(x_{i},x_{j},i-j)$ such that
@@
where $f_{q}$ and $f_{k}$ denote the neural network embeddings of query and key inside an attention layer. Fortunately, there is an explicit way of building $g$ by multiplying $f_{\{q,k\}}(x_{i},i)$ with the diagonal matrix expanded by rotation matrices (Su et al., 2024):
@@
Temporal connection. In addition to commonly adapted message passing in the spatial domain, we consider another dimension, temporal domain, which can be described by the depth of the neural network (Chen et al., 2018). As message-passing neural networks are local, we believe it is beneficial to integrate multi-scale information. Specifically, for each consecutive layer, we learn a kernel that linear transforms and aggregates features. Given the invariant feature embeddings $h^{l_{1}}$ at layer $l_{1}$ and $h^{(l_{2})}$ at layer $l_{2}$, the kernel is of the matrix product form: $M^{(l_{3,1,2})}$ which transforms $h^{(l_{1})}$ and $h^{(l_{2})}$ to a new feature $h^{(l_{3})}$ and added as a residual into the next layer:
@@
Supplementary Information
AlphaNet: Scaling Up Local Frame-based NNIP
@@ Appendix A Local Structures, Completeness, and Expressive Power of GNNs
The expressive power of permutation-invariant GNNs is classically evaluated via the 1-Weisfeiler-Lehman (1-WL) test  (Xu et al., 2019). As established by  (Wijesinghe and Wang, 2022), the 1-WL test is equivalent to a GNN’s capacity to discriminate local subtree isomorphism, measuring its ability to distinguish hierarchical neighborhood structures. However, this framework operates solely on 2D graph topologies, ignoring 3D geometric relationships (e.g., bond angles, torsional dihedrals), thereby failing to address molecular or material systems where spatial configurations are decisive.
Completeness in GNNs requires a transformation function $f$ to satisfy $f(X)=f(Y)\iff X\text{ and }Y\text{ are isomorphic}$. We can extends completeness to local geometric isomorphism, where isomorphism is defined under 3D geometric constraints (distances, angles, chiralities). Concurrently, methods encoding 3D geometric invariants  (Wijesinghe and Wang, 2022) have demonstrated enhanced expressivity, enabling differentiation of graphs with distinct local atomic environments beyond WL limitations. Similar to LeftNet (Du et al., 2023), AlphaNet is able to differentiate local structures beyond tree isomorphism by defining the 3D structure weights.
@@ B.1 Datasets
The Defected Bilayer Graphene Dataset (Ying et al., 2024) consists of reference structures specifically designed to train and validate machine learning potentials (MLPs). It includes three bilayer systems: V0V0 (pristine), V0V1 (single vacancy on the top layer), and V1V1 (single vacancy in both layers). The dataset comprises single-point DFT (PBE+MBD) energies and atomic forces calculated for various configurations, including different interlayer distances, stacking modes, and manually deformed structures. Additionally, snapshot configurations from MD simulations at different temperatures were included. The data were split into training, validation, and test sets, containing 3988, 4467, and 200 structures, respectively. Farthest point sampling (FPS) and principal component analysis (PCA) were used to ensure a representative and balanced division.
@@
The Formate Decomposition on Cu Dataset (Batzner et al., 2022) consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond. The dataset includes initial, intermediate, and final states, such as monodentate and bidentate formate on Cu <110>, as well as the final state with an H ad-atom and desorbed CO2 in the gas phase. Nudged elastic band (NEB) method (Henkelman et al., 2000) was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations. These simulations were performed using the CP2K code, resulting in a total of 6855 DFT structures with a time step of 0.5 fs and 500 steps per trajectory. AlphaNet was trained on 2500 uniformly sampled structures from the full dataset, with a validation set of 250 structures and the mean absolute error evaluated on the remaining structures.
@@
The Zeolite Dataset comprises essential reference structures designed for various applications such as catalysis, adsorption, and separation. Zeolites are highly valued porous materials that have been the subject of extensive scientific research due to their broad applicability. The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) (Car and Parrinello, 1985) simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) (Kresse and Furthmüller, 1996; Kresse and Hafner, 1993). For each zeolite, 80,000 snapshots were extracted, and the corresponding energies and atomic forces were calculated for every configuration. The dataset was randomly divided into training, validation, and test sets in a 6:2:2 ratio, comprising 48,000, 16,000, and 16,000 structures, respectively.
@@
The OC2M Dataset (Chanussot et al., 2021) is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials. This dataset includes approximately 2 million atomic configurations in the training set, with validation and test sets containing 4 splits, which is In Domain (ID), out-of-domain adsorbate (OOD adsorbate), out-of-domain catalyst (OOD catalyst), and both of unseen adsorbate and unseen catalyst (OOD both). Each of them contains around 1 million structures. It encompasses configurations involving 56 different chemical elements, ensuring a broad and diverse representation of atomic environments. These configurations were generated using density functional theory (DFT) calculations, with a focus on a wide range of materials science applications, including bulk materials, surfaces, and defect structures. Originating from the comprehensive OC20 Dataset, the OC2M subset provides accurate total energies and atomic forces for each configuration. This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
@@
The MPTraj Dataset and Matbench Discovery Test (Riebesell et al., 2024). MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials (Deng et al., 2023). Matbench Discovery is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set (Wang et al., 2021). In this case we train our model on the MPTraj dataset and test it on the WBM test set.
@@ B.4 Baseline Models
For computational modeling configurations, NequIP achieves optimal performance with tensor rank 2, feature size 32, and 5 Åcutoff in the formate decomposition dataset, while using tensor rank 1 with 7 Åcutoff for defective graphene studies. The zeolite dataset employs a fine-tuned DPA-2 potential via Deep Pot implementation ((Zhang et al., 2024b)). Models trained on OC20 leverage pre-trained SchNet/DimeNet++ architectures from https://fair-chem.github.io/core/model_checkpoints.html, with other results cross-validated against (Qu and Krishnapriyan, 2024) using publicly accessible validation sets. MPtrj benchmark data integrates results from https://matbench-discovery.materialsproject.org/ and comparative analysis in (Qu and Krishnapriyan, 2024). Speed testing on zeolite structures maintains uniform 2-layer architectures with 5 Åcutoff, where PaiNN/SchNet employ 120-atom basis and 16 RBF basis, MACE uses tensor order 0-1 (feature size 16), NequIP extends to tensor order 0-2 (feature size 16), AlphaNet configures 32 hidden channels with 12 attention heads, and EquiformerV2 implements sphere/edge channels 16 with 16-dimensional attention hidden states and 2-head mechanisms.
@@ Appendix D Data Scaling and Dynamic loss weights
The experimental datasets were generated using VASP (Kresse and Furthmüller, 1996; Kresse and Hafner, 1993), where the energy per atom typically converges around $5\,\mathrm{eV}$ under our default training settings. These configurations demonstrated robust performance within this energy regime. However, when applied to systems with significantly higher energy magnitudes per atom and greater energy variance—such as those produced by CP2K (Kühne et al., 2020) simulations—the original settings failed to achieve stable convergence. To address this challenge, we developed a dynamic optimization strategy: initially adjusting the energy-to-force loss weight ratio from $\lambda_{E}:\lambda_{F}=4:100$ to $0.05:100$, then progressively increasing $\lambda_{E}$ while simultaneously decaying the learning rate. This coordinated adaptation proved critical for successful training in high-variance scenarios.

### v4

@@ 2.3 AlphaNet demonstrating enhanced speed-accuracy trade-off
The final efficiency evaluation is conducted using the OC20 dataset, with results summarized in Table 5. Notably, EScAIP demonstrates outstanding efficiency in both inference speed and memory usage, even given its substantial model size. AlphaNet exhibits a marginally faster inference speed; however, it incurs higher memory consumption compared to EScAIP. It is important to emphasize that models such as AlphaNet, SchNet, and DimeNet++ are conservative in their force calculations, inherently relying on PyTorch’s autograd functionality, thus necessitating additional computational overhead and consequently increasing inference time and memory usage.
@@ 3 Discussion
AlphaNet represents a significant advance in scalable neural network interatomic potentials, achieving exceptional accuracy while maintaining computational efficiency. Its success stems from a novel local-frame-based architecture that scalarizes geometric information within local frames and effectively aggregates these frames to produce features sensitive to both local and global chemical environments. Leveraging this architecture, AlphaNet adeptly captures complex atomic interactions across diverse chemical bonding types, including metallic, covalent, ionic, and long-range interactions, demonstrating versatility and broad applicability.
One of AlphaNet’s key strengths is its exceptional scalability. The model consistently maintains high performance across varying dataset size and complexity of atomic systems. Remarkably, AlphaNet achieves robust accuracy while utilizing fewer parameters compared to conventional approaches, enhancing its computational efficiency. This capability to effectively manage large-scale datasets without compromising performance positions AlphaNet as a highly promising framework for extensive simulations and practical real-world applications.
In future work, we plan to extend the model’s capabilities to account for additional interaction types, notably hydrogen bonding, which plays a critical role across diverse chemical and biological systems. Another important direction is enhancing the interpretability of the proposed model. Although our current implementation demonstrates strong predictive performance, its applicability and reliability in practical scenarios can be significantly improved by achieving deeper insight into how the model identifies, processes, and leverages atomic-level interactions.
Additionally, addressing the efficient use of GPU memory during training presents a long-term challenge, particularly for large systems exceeding 1,000 atoms—even when leveraging parallelization across multiple GPUs. This limitation, common among most message-passing neural networks, must be overcome to facilitate scalability toward larger and more complex systems. Future research will prioritize memory optimization strategies to surmount this bottleneck, ensuring broader applicability and enabling analysis of increasingly intricate molecular systems.
@@ 4 Methodology
Equivariant message-passing neural networks. Equivariance is another crucial property to building MLFFs as molecular systems are invariant to the Euclidean group. A function $f$ is said to be equivariant with respect to certain group $G$ if $f\circ g(x)=g\circ f(x)$, $\forall x\in X,\,g\in G$. One natural property of message-passing neural networks is the permutation equivariance such that changing the order of the input set of nodes do not result in difference output for the same node as it only depends on the neighbors which are unchanged under permutation. In addition to permutation, we need to consider the special Euclidean group in 3D, SE(3), which includes rotation and translation operations. We do so by leveraging local frames [16]. Specifically, we build a set of equivariant and complete frames and scalarize the geometric quantity which does not break SE(3)-equivariance under nonlinear neural network encodings.
To build equivariant and complete frames from atomic positions, we consider the following edge-based frames following [16]:
@@
Efficient and expressive equivariant message-passing neural networks. Following [18], we leverage an efficient and expressive extension of the above equivariant message-passing neural networks based on local frames. We consider two modules to improve the expressiveness while maintaining the efficiency of the network. Specifically, we leverage a local structure encoding module which scalarizes not only features from local neighbors but also from overlapping neighbors along each edge, denoted as $A_{ij}:=f_{l}(\text{Scalarize}(S_{i-j},\mathcal{F}_{ij}))$, where $f_{l}$ denotes a neural network and $S_{i-j}$ represents overlapping neighbors of node $i$ and $j$. In addition, we use an additional frame transition module which encodes the alignment between each neighboring frame to improve the expressiveness. We realize this by incorporating vector-based message passing to amortize the cost where $\mathbb{m}_{ij}$ denotes equivariant edge features. This framework is highly efficient as it avoids expensive higher-body message-passing neural networks and higher-order tensor updates.
@@
Rotary positional embedding as invariant frame transition. Positional embedding are known to improve and stabilize the training of transformers [48]. We introduce the elegant relationship between the commonly used Rotary Position Embedding (RoPE) and frame transition we use in our architecture [52]. In fact the frame transition can be considered as a generalized rotary positional embedding for 3D equivariant frames. Note that the spirit of rotary position embedding is to design an embedding scheme for relative position in sequential data. To build the rotary positional embedding,  [52] utilizes the relative index between nodes $x_{i}$ and $x_{j}$ of a one-dimensional sequence to find a function $g(x_{i},x_{j},i-j)$ such that
@@
where $f_{q}$ and $f_{k}$ denote the neural network embeddings of query and key inside an attention layer. Fortunately, there is an explicit way of building $g$ by multiplying $f_{\{q,k\}}(x_{i},i)$ with the diagonal matrix expanded by rotation matrices [52]:
@@
Temporal connection. In addition to commonly adapted message passing in the spatial domain, we consider another dimension, temporal domain, which can be described by the depth of the neural network [12]. As message-passing neural networks are local, we believe it is beneficial to integrate multi-scale information. Specifically, for each consecutive layer, we learn a kernel that linear transforms and aggregates features. Given the invariant feature embeddings $h^{l_{1}}$ at layer $l_{1}$ and $h^{(l_{2})}$ at layer $l_{2}$, the kernel is of the matrix product form: $M^{(l_{3,1,2})}$ which transforms $h^{(l_{1})}$ and $h^{(l_{2})}$ to a new feature $h^{(l_{3})}$ and added as a residual into the next layer:
@@
Supplementary Information for
AlphaNet: Scaling Up Local-frame-based Neural Network Interatomic Potentials
@@ Appendix A Local Structures, Completeness, and Expressive Power of GNNs
The expressive power of permutation-invariant GNNs is classically evaluated via the 1-Weisfeiler-Lehman (1-WL) test  [59]. As established by  [58], the 1-WL test is equivalent to a GNN’s capacity to discriminate local subtree isomorphism, measuring its ability to distinguish hierarchical neighborhood structures. However, this framework operates solely on 2D graph topologies, ignoring 3D geometric relationships (e.g., bond angles, torsional dihedrals), thereby failing to address molecular or material systems where spatial configurations are decisive.
Completeness in GNNs requires a transformation function $f$ to satisfy $f(X)=f(Y)\iff X\text{ and }Y\text{ are isomorphic}$. We can extends completeness to local geometric isomorphism, where isomorphism is defined under 3D geometric constraints (distances, angles, chiralities). Concurrently, methods encoding 3D geometric invariants  [58] have demonstrated enhanced expressivity, enabling differentiation of graphs with distinct local atomic environments beyond WL limitations. Similar to LeftNet [17], AlphaNet is able to differentiate local structures beyond tree isomorphism by defining the 3D structure weights.
@@ B.1 Datasets
The Defected Bilayer Graphene Dataset [61] consists of reference structures specifically designed to train and validate machine learning potentials (MLPs). It includes three bilayer systems: V0V0 (pristine), V0V1 (single vacancy on the top layer), and V1V1 (single vacancy in both layers). The dataset comprises single-point DFT (PBE+MBD) energies and atomic forces calculated for various configurations, including different interlayer distances, stacking modes, and manually deformed structures. Additionally, snapshot configurations from MD simulations at different temperatures were included. The data were split into training, validation, and test sets, containing 3988, 4467, and 200 structures, respectively. Farthest point sampling (FPS) and principal component analysis (PCA) were used to ensure a representative and balanced division.
@@
The Formate Decomposition on Cu Dataset [5] consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond. The dataset includes initial, intermediate, and final states, such as monodentate and bidentate formate on Cu <110>, as well as the final state with an H ad-atom and desorbed CO2 in the gas phase. Nudged elastic band (NEB) method [26] was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations. These simulations were performed using the CP2K code, resulting in a total of 6855 DFT structures with a time step of 0.5 fs and 500 steps per trajectory. AlphaNet was trained on 2500 uniformly sampled structures from the full dataset, with a validation set of 250 structures and the mean absolute error evaluated on the remaining structures.
@@
The Zeolite Dataset comprises essential reference structures designed for various applications such as catalysis, adsorption, and separation. Zeolites are highly valued porous materials that have been the subject of extensive scientific research due to their broad applicability. The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) [10] simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) [33, 34]. For each zeolite, 80,000 snapshots were extracted, and the corresponding energies and atomic forces were calculated for every configuration. The dataset was randomly divided into training, validation, and test sets in a 6:2:2 ratio, comprising 48,000, 16,000, and 16,000 structures, respectively.
@@
The OC2M Dataset [11] is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials. This dataset includes approximately 2 million atomic configurations in the training set, with validation and test sets containing 4 splits, which is In Domain (ID), out-of-domain adsorbate (OOD adsorbate), out-of-domain catalyst (OOD catalyst), and both of unseen adsorbate and unseen catalyst (OOD both). Each of them contains around 1 million structures. It encompasses configurations involving 56 different chemical elements, ensuring a broad and diverse representation of atomic environments. These configurations were generated using density functional theory (DFT) calculations, with a focus on a wide range of materials science applications, including bulk materials, surfaces, and defect structures. Originating from the comprehensive OC20 Dataset, the OC2M subset provides accurate total energies and atomic forces for each configuration. This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
@@
The MPTraj Dataset and Matbench Discovery Test [44]. MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials [15]. Matbench Discovery is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set [57]. In this case we train our model on the MPTraj dataset and test it on the WBM test set.
@@ B.4 Baseline Models
For computational modeling configurations, NequIP achieves optimal performance with tensor rank 2, feature size 32, and 5 Åcutoff in the formate decomposition dataset, while using tensor rank 1 with 7 Åcutoff for defective graphene studies. The zeolite dataset employs a fine-tuned DPA-2 potential via Deep Pot implementation ([63]). Models trained on OC20 leverage pre-trained SchNet/DimeNet++ architectures from https://fair-chem.github.io/core/model_checkpoints.html, with other results cross-validated against [42] using publicly accessible validation sets. MPtrj benchmark data integrates results from https://matbench-discovery.materialsproject.org/ and comparative analysis in [42]. Speed testing on zeolite structures maintains uniform 2-layer architectures with 5 Åcutoff, where PaiNN/SchNet employ 120-atom basis and 16 RBF basis, MACE uses tensor order 0-1 (feature size 16), NequIP extends to tensor order 0-2 (feature size 16), AlphaNet configures 32 hidden channels with 12 attention heads, and EquiformerV2 implements sphere/edge channels 16 with 16-dimensional attention hidden states and 2-head mechanisms.
@@ Appendix D Data Scaling and Dynamic loss weights
The experimental datasets were generated using VASP [33, 34], where the energy per atom typically converges around $5\,\mathrm{eV}$ under our default training settings. These configurations demonstrated robust performance within this energy regime. However, when applied to systems with significantly higher energy magnitudes per atom and greater energy variance—such as those produced by CP2K [35] simulations—the original settings failed to achieve stable convergence. To address this challenge, we developed a dynamic optimization strategy: initially adjusting the energy-to-force loss weight ratio from $\lambda_{E}:\lambda_{F}=4:100$ to $0.05:100$, then progressively increasing $\lambda_{E}$ while simultaneously decaying the learning rate. This coordinated adaptation proved critical for successful training in high-variance scenarios.